# Text-controlled Motion Mamba: Text-Instructed Temporal Grounding of Human Motion

### v1

@@ 5.4. More Analysis
The memory usage of the Mamba-based models increases linearly with the length of the sequence, which allows them to effectively handle longer sequences. Figure 4 validates this by comparing the GPU memory consumption of our TM-Mamba model with its transformer and LSTM counterparts when processing sequences of varying lengths. As depicted in the figure, the memory cost of the transformer-based model increases quadratically with sequence length, causing it to quickly run out of memory. On the other hand, both the LSTM and TM-Mamba models exhibit linear memory consumption, making them applicable for processing long motion sequences when extracting global context information. The slope of TM-Mamba is larger than its LSTM counterpart due to larger hidden states.
@@
Table 3 presents a comparison between TM-Mamba and its temporal transformer counterpart across various maximum sequence lengths on the BABEL-Grounding dataset. As shown, TM-Mamba exhibits superior performance to temporal transformer on shorter motion sequence. Meanwhile, the transformer model runs out of GPU memory at sequence lengths exceeding 1000 due to its quadratic memory consumption. In contrast, TM-Mamba, benefiting from its linear memory cost, manages longer sequences effectively. It can be observed that increasing sequence length complicates the extraction of temporal global context, leading to a decline in motion grounding performance.
@@
In order to demonstrate the effectiveness of our proposed model, we provide visualizations on the BABEL-Grounding dataset to showcase the comparisons between the TM-Mamba model and two ablative models (w/o text control and w/o relational embedding). Figure 5 presents the frame activation scores and the ground-truth temporal segments corresponding to the text query. As illustrated, our full model manifests higher activation scores within the ground-truth temporal segments in contrast to other models, thereby resulting in enhanced accuracy in the grounding performance. Figure 6 further visualizes the grounding results predicted by the model in comparison with the ground-truth. As depicted, the ablative models struggle to accurately predict the time spans corresponding to the text query, while TM-Mamba achieves significantly improved grounding precision.
6. Conclusion
This work introduces a novel task called text-based human motion grounding (THMG), which aims to determine the start and end timestamps of all segments from an untrimmed motion sequence given a textual description. The key challenge lies in extracting global temporal information from lengthy untrimmed sequences based on text query, while transformer-based methods suffer from quadratic memory cost. To this end, we draw inspiration from recent advances in state space models, and propose a unified framework called TM-Mamba with linear memory cost. TM-Mamba incorporates a novel text-controlled selection mechanism into the Mamba algorithm, enabling the model to dynamically propagate input information based on text queries and extract relevant global context. A relational embedding is incorporated to model the underlying graph topology of the human skeleton. For evaluation, a text-motion dataset called BABEL-Grounding is constructed, which is the first one that provides detailed textual descriptions with their corresponding temporal segments annotation. Rigorous experiments demonstrate the effectiveness of the proposed model.

### v2

@@ V-D More Analysis
The memory usage of the Mamba-based models increases linearly with the length of the sequence, which allows them to effectively handle longer sequences. Figure 8 validates this by comparing the GPU memory consumption of our TM-Mamba model with its Transformer, LSTM and flash-attention [128] counterparts when processing sequences of varying lengths. As depicted in the figure, the memory cost of the Transformer-based model increases quadratically with sequence length, causing it to quickly run out of memory. On the other hand, TM-Mamba, LSTM and flash-attention-based efficient transformer exhibit linear memory consumption, making them applicable for processing long motion sequences when extracting global context information. The slope of TM-Mamba is larger than its LSTM counterpart due to larger hidden states.
@@
Table IV presents a comparison between TM-Mamba and its temporal Transformer counterpart across various maximum sequence lengths on the BABEL-Grounding dataset. Specifically, the Transformer baseline consists of a transformer encoder for encoding the motion sequence and a cross-attention module for injecting textual conditions. In our experiments, we configure the model with the following parameters: $d_{model}=256$, $d_{feedforward}=2048$, and $nhead=8$. The motion encoder contains 6 layers, while the cross-attention module contains 3 layers. Here, the feature dimension $d_{model}=256$ is the same as in TM-Mamba. As shown, TM-Mamba exhibits superior performance to temporal Transformer on shorter motion sequence. Meanwhile, the Transformer model runs out of GPU memory at sequence lengths exceeding 1000 due to its quadratic memory consumption. In contrast, TM-Mamba, benefiting from its linear memory cost, manages longer sequences effectively. It can be observed that increasing sequence length complicates the extraction of temporal global context, leading to a decline in motion grounding performance. To ensure the performance gains are not due to more parameters, we compared the models’ parameter counts. The number of parameters for each model is as follows (the CLIP text encoder clip-vit-base-patch32 is excluded from the parameter count):
  • Transformer: 13.72M parameters
  • TM-Mamba: 6.25M parameters
Therefore, the performance improvement cannot be attributed to a larger parameter size.
TABLE V: Impact of the number of TM-Mamba blocks on the grounding performance. GPU memory overflow occurred with 4 blocks.
Blocks mAP@IoU (%)
0.1 0.2 0.3 0.4 0.5 0.6 0.7 Avg
1 47.8 44.1 40.3 36.4 32.9 28.2 22.8 36.1
2 53.6 50.0 46.0 42.0 37.5 32.0 25.5 40.9
3 53.9 50.5 46.7 42.8 38.4 32.6 26.0 41.6
Figure 11: Visualizations of the grounding results of TM-Mamba compared to baseline models. ‘GT’ denotes ground-truth temporal segments corresponding to the text query. Best viewed in color.
@@
In order to demonstrate the effectiveness of our proposed model, we provide a series of visualizations on the BABEL-Grounding dataset. Figure 9 presents the frame activation scores and the ground-truth temporal segments corresponding to the text query to showcase the comparisons between the TM-Mamba and two ablative models (w/o text control and w/o relational embedding). As illustrated, our full model manifests higher activation scores within the ground-truth temporal segments in contrast to other models, thereby resulting in enhanced accuracy in the grounding performance. Figure 10 and Figure 11 further visualizes the grounding results predicted by TM-Mamba in comparison with ablative and baseline models respectively. As depicted, the comparative models struggle to accurately predict the time spans corresponding to the text query, while TM-Mamba achieves significantly improved grounding precision.
Number of Blocks
Table V illustrates the influence of the number of TM-Mamba blocks on grounding performance. Notably, performance begins to saturate when the number reaches 3 blocks, and encounter a GPU memory overflow at 4 blocks. We set the number of blocks as 3 to strike a balance between performance and computational efficiency.
VI Conclusion
This work introduces a novel task called text-based human motion grounding (THMG), which aims to determine the start and end timestamps of all segments from an untrimmed motion sequence given a textual description. The key challenge lies in extracting global temporal information from lengthy untrimmed sequences based on text query, while Transformer-based methods suffer from quadratic memory cost. To this end, we draw inspiration from recent advances in state space models, and propose a unified framework called TM-Mamba with linear memory cost. TM-Mamba incorporates a novel text-controlled selection mechanism into the Mamba algorithm, enabling the model to dynamically propagate input information based on text queries and extract relevant global context. A relational embedding is incorporated to model the graph topology of the human skeleton. For evaluation, a text-motion dataset called BABEL-Grounding is constructed, which is the first one that provides detailed textual descriptions with their corresponding temporal segments annotation. Rigorous experiments validate the effectiveness of the proposed model.